# Distributional Reduction: Unifying Dimensionality Reduction and Clustering with Gromov-Wasserstein

### v2

@@ 5 Numerical Experiments
In this section, we demonstrate the relevance of our approach for joint clustering and DR, over 8 labeled datasets detailed in Appendix E including: 3 image datasets (COIL-20 nene1996columbia , MNIST & fashion-MNIST xiao2017fashion ) and 5 genomic ones (PBMC wolf2018scanpy , SNA 1 & 2 chen2019high and ZEISEL 1 & 2zeisel2015cell ). To this end, we examine how effectively DistR prototypes offer discriminative representations of input samples at various granularities, using an evaluation system described below.
@@
Evaluation metrics. We consider several evaluation metrics to quantify whether the output prototypes provide a meaningful summarized representation of the input datasets. As often done in the literature, we focus on supervised evaluations while assessing two main properties. Firstly, we wonder whether the prototypes are individually discriminating input samples $\{{\mathbf{x}}_{i}\}$ w.r.t their class labels $\{y_{i}\}$ via their associated prototype. We thus compute the homogeneity score rosenberg2007v that quantifies to which extent each prototype ${\mathbf{z}}_{k}$ is associated with points in the same class. Secondly, we aim to evaluate whether the embedding space is both globally discriminant, i.e. that prototypes organize themselves to form clusters that match the classes, measured by two metrics. i) We associate with each prototype ${\mathbf{z}}_{k}$ a label using majority voting from the associated samples. Then we make use of the popular silhouette score rousseeuw1987silhouettes to evaluate whether prototypes’ positions are correctly aligned with these labels. ii) As commonly done to assess DR representations huang2022towards , we also consider the downstream task of clustering the prototypes using the k-means algorithm. We then assign to each $\{{\mathbf{x}}_{i}\}$ the cluster of its associated prototype, acting as a predicted label, and compute the Normalized Mutual Information (NMI) kvaalseth2017normalized between these labels and $\{y_{i}\}$. More details about evaluation metrics can be found in Appendix E.
@@
Figure 4: Scores (×$\times 100$) with respect to the number of prototypes (in $\mathbb{R}^{2}$) produced by DistR using the t-SNE model: SEA similarity for ${\mathbf{C}}_{X}$ van2023snekhorn , Student’s kernel for ${\mathbf{C}}_{Z}$ and loss $L_{\operatorname{KL}}$.
Figure 5: Examples of 2-dimensional embeddings produced by DistR using the SEA similarity for ${\mathbf{C}}_{X}$ and the Student’s kernel for ${\mathbf{C}}_{Z}$. The latter computed in $\mathbb{R}^{2}$ for the first three datasets and the Poincaré ball for the last one. Displayed images are medoids for each cluster i.e. $\operatorname*{arg\,max}_{i}[{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}_{:,k}]_{i}$ for cluster $k$. The area of image $k$ is proportional to $[{\mathbf{h}}_{Z}]_{k}$.
@@ 6 Conclusion
By making a connection between the GW problem and popular clustering and DR algorithms, we proposed a unifying framework denoted as DistR. DistR enables transforming any DR algorithm into a joint DR-clustering method that produces embeddings (or prototypes) with chosen granularity. We believe that the versatility of the GW framework will enable new extensions in unsupervised learning. For instance, the formalism associated with (semi-relaxed) GW barycenters naturally enables considering multiple inputs, potentially unaligned and of different sizes. Hence our approach can be useful to address challenges related to the multi-view dimensionality reduction and clustering problems. Other promising directions involve, better capturing multiple dependency scales in the input data by hierarchically adapting the resolution of the embedding similarity graph, or enabling batch optimization of embeddings to operate over much larger datasets.
Appendix A Proof of results in Section 3
A.1 Proof of lemma 3.1
@@ Appendix A Proof of results in Section 3
A.2 Proof of theorem 3.2
In the following $\operatorname{DS}$ is the space of $N\times N$ doubly stochastic matrices. We begin by proving the first point of theorem 3.2. We will rely on the simple, but useful, result below
Proposition A.1.
@@
$\begin{split}\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij})&=\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},\sum_{kl}[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{kl}T_{ik}T_{jl})\\ &\leq\sum_{ijkl}L([{\mathbf{C}}_{\mathbf{X}}]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{kl})T_{ik}T_{jl}\,.\end{split}$ (7)
@@
$\min_{{\mathbf{Z}}}\min_{{\mathbf{T}}\in\operatorname{DS}}\ \sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij})\leq\min_{{\mathbf{Z}}}\ \min_{{\mathbf{T}}\in\operatorname{DS}}\sum_{ijkl}L([{\mathbf{C}}_{\mathbf{X}}]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{kl})T_{ik}T_{jl}\,.$ (8)
@@
For the first point it suffices to see that the condition eq. 11 implies that $\{{\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}:{\mathbf{Z}}\in\mathbb{R}^{N\times d},{\mathbf{T}}\in\operatorname{DS}\}\subseteq\{{\mathbf{C}}_{Z}({\mathbf{Z}}):{\mathbf{Z}}\in\mathbb{R}^{N\times d}\}$ (in fact we have equality by choosing ${\mathbf{T}}=\mathbf{I}_{N}$) and thus eq. 6 holds and we apply proposition A.1.
For the second point we will also show that eq. 6 holds. Consider ${\mathbf{Z}}^{\star},{\mathbf{T}}^{\star}$ minimizers of $\min_{{\mathbf{Z}},{\mathbf{T}}\in\operatorname{DS}}\ \sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij})$. By hypothesis there exists ${\mathbf{Y}}\in\mathbb{R}^{N\times d}$ such that $\forall(i,j)\in{[\![N]\!]}^{2},[{\mathbf{T}}^{\star}{\mathbf{C}}_{Z}({\mathbf{Z}}^{\star}){{\mathbf{T}}^{\star}}^{\top}]_{ij}\geq[{\mathbf{C}}_{Z}({\mathbf{Y}})]_{ij}$. Since $L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},\cdot)$ is non-decreasing for any $(i,j)$ then $\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}^{\star}{\mathbf{C}}_{Z}({\mathbf{Z}}^{\star}){{\mathbf{T}}^{\star}}^{\top}]_{ij})\geq\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Y}})]_{ij})$ and thus $\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}^{\star}{\mathbf{C}}_{Z}({\mathbf{Z}}^{\star}){{\mathbf{T}}^{\star}}^{\top}]_{ij})\geq\min_{{\mathbf{Z}}}\ \sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij})$ which gives the condition eq. 6 and we have the conclusion by proposition A.1. ∎
@@
    $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\sum_{ij}L_{2}([{\mathbf{C}}_{\mathbf{X}}]_{ij},\langle{\mathbf{z}}_{i},{\mathbf{z}}_{j}\rangle)$ (13)
@@
For the first point we show that the condition eq. 11 of proposition A.2 is satisfied. Indeed take any ${\mathbf{Z}},{\mathbf{T}}$ then ${\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}={\mathbf{T}}{\mathbf{Z}}{\mathbf{Z}}^{\top}{\mathbf{T}}^{\top}=({\mathbf{T}}{\mathbf{Z}})({\mathbf{T}}{\mathbf{Z}})^{\top}={\mathbf{C}}_{Z}({\mathbf{T}}{\mathbf{Z}})$.
For the second point if we consider $\tilde{L}(a,b)=a\times b$ then we use that the neighbor embedding problem eq. 16 is equivalent to $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\sum_{ij}-[{\mathbf{C}}_{X}({\mathbf{X}})]_{ij}\log([{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij})=\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\sum_{ij}\tilde{L}([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}})]_{ij})$ where $[\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}})]_{ij}=g({\mathbf{z}}_{i}-{\mathbf{z}}_{j})+\log(R({\mathbf{Z}}))$ with $g=-\log\circ f$. Since $f$ is logarithmically concave $g$ is convex. Moreover we have that $\tilde{L}(a,\cdot)$ is convex (it is linear) and non-decreasing since $a\in\mathbb{R}_{+}$ in this case (${\mathbf{C}}_{\mathbf{X}}$ is non-negative). Also for any ${\mathbf{Z}}\in\mathbb{R}^{N\times d}$ and ${\mathbf{T}}\in\operatorname{DS}$ we have, using Jensen’s inequality since ${\mathbf{T}}$ is doubly-stochastic,
@@
Now we will prove that $\log(R({\mathbf{Z}}))\geq\log(R({\mathbf{T}}{\mathbf{Z}}))$. Using Birkhoff’s theorem [6] the matrix ${\mathbf{T}}$ can be decomposed as a convex combination of permutation matrices, i.e. , ${\mathbf{T}}=\sum_{k}\lambda_{k}{\mathbf{P}}_{k}$ where $({\mathbf{P}}_{k})_{k}$ are permutation matrices and $\lambda_{k}\in\mathbb{R}_{+}$ with $\sum_{k}\lambda_{k}=1$. Hence by convexity and Jensen’s inequality $R({\mathbf{T}}{\mathbf{Z}})=R(\sum_{k}\lambda_{k}{\mathbf{P}}_{k}{\mathbf{Z}})\leq\sum_{k}\lambda_{k}R({\mathbf{P}}_{k}{\mathbf{Z}})$. Now using that $R$ is permutation invariant we get $R({\mathbf{P}}_{k}{\mathbf{Z}})=R({\mathbf{Z}})$ and thus $R({\mathbf{T}}{\mathbf{Z}})\leq\sum_{k}\lambda_{k}R({\mathbf{Z}})=R({\mathbf{Z}})$. Since the logarithm is non-decreasing we have $\log(R({\mathbf{Z}}))\geq\log(R({\mathbf{T}}{\mathbf{Z}}))$ and, overall,
@@
Thus if we introduce ${\mathbf{Y}}={\mathbf{T}}{\mathbf{Z}}$ we have $[{\mathbf{T}}\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij}\geq[\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Y}})]_{ij}$ and eq. 12 is satisfied. Thus we can apply proposition A.2 and state that $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\sum_{ij}\tilde{L}([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}})]_{ij})$ and $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\operatorname{GW}_{\tilde{L}}({\mathbf{C}}_{X}({\mathbf{X}}),\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}}),\frac{1}{N}\mathbf{1}_{N},\frac{1}{N}\mathbf{1}_{N})$ are equivalent which concludes as $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\operatorname{GW}_{\tilde{L}}({\mathbf{C}}_{X}({\mathbf{X}}),\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}}),\frac{1}{N}\mathbf{1}_{N},\frac{1}{N}\mathbf{1}_{N})=\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),\frac{1}{N}\mathbf{1}_{N},\frac{1}{N}\mathbf{1}_{N})$. ∎
@@
is actually concave. Indeed, in this case there exists a minimizer which is an extremal point of ${\mathcal{U}}(\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$. By Birkhoff’s theorem [6] these extreme points are the matrices $\frac{1}{N}{\mathbf{P}}$ where ${\mathbf{P}}$ is a $N\times N$ permutation matrix. Consequently, when the function eq. 21 is concave minimizing $\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$ in ${\mathbf{Z}}$ is equivalent to minimizing in ${\mathbf{Z}}$
@@
where $a_{\mathbf{X}},b_{\mathbf{Z}}$ are terms that do not depend on ${\mathbf{T}}$. Since the problem is quadratic the concavity only depends on the term $-\sum_{ijkl}[{\mathbf{C}}_{X}({\mathbf{X}})]_{ik}\log([{\mathbf{K}}_{\mathbf{Z}}]_{jl})T_{ij}T_{kl}=-\operatorname{Tr}({\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}^{\top}\log({\mathbf{K}}_{\mathbf{Z}}){\mathbf{T}})$. From [47] we know that the function ${\mathbf{T}}\to-\operatorname{Tr}({\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}^{\top}\log({\mathbf{K}}_{\mathbf{Z}}){\mathbf{T}})$ is concave on ${\mathcal{U}}(\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$ when ${\mathbf{C}}_{X}({\mathbf{X}})$ is CPD and $\log({\mathbf{W}}_{\mathbf{Z}})$ is CPD. This concludes the proof.
A.3 Necessary and sufficient condition

### v3

@@ 5 Numerical Experiments
In this section, we demonstrate the relevance of our approach for joint clustering and DR, over 8 labeled datasets detailed in Appendix F including: 3 image datasets (COIL-20 Nene et al. 1996, MNIST & fashion-MNIST Xiao et al. 2017) and 5 genomic ones (PBMC Wolf et al. 2018, SNA 1 & 2 Chen et al. 2019 and ZEISEL 1 & 2Zeisel et al. 2015). To this end, we examine how effectively DistR prototypes offer discriminative representations of input samples at various granularities, using an evaluation system described below. Code is provided at https://github.com/huguesva/Distributional-Reduction.
@@
Figure 3: Best trade-off between homogeneity vs silhouette (2 first plots), and homogeneity vs NMI (2 last plots). Scores are normalized in $\left[0,1\right]$ via min-max scaling over a dataset. Small markers are scores per dataset for 5 runs, while big ones are their mean. We illustrate the 20-80% percentiles of scores per method as a colored surface.
Figure 4: Examples of embeddings produced by DistR, using the SEA similarity for ${\mathbf{C}}_{X}$ and the Student’s kernel for ${\mathbf{C}}_{Z}$, respectively in $\mathbb{R}^{2}$ for the first three datasets and the Poincaré ball for the last one. Displayed images are medoids for each cluster i.e. $\operatorname*{arg\,max}_{i}[{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}_{:,k}]_{i}$ for cluster $k$. The area of image $k$ is proportional to $[{\mathbf{h}}_{Z}]_{k}$.
Figure 5: Scores (×$\times 100$) w.r.t with respect to the number of prototypes (in $\mathbb{R}^{2}$) using the t-SNE model: SEA similarity for ${\mathbf{C}}_{X}$ (Van Assel et al., 2023), Student’s kernel for ${\mathbf{C}}_{Z}$ and loss $L_{\operatorname{KL}}$.
Evaluation metrics. We consider several evaluation metrics. Firstly, we wonder whether the prototypes are individually discriminating input samples $\{{\mathbf{x}}_{i}\}$ w.r.t their class label. We consider the homogeneity score (Rosenberg and Hirschberg, 2007) that quantifies to which extent each prototype ${\mathbf{z}}_{k}$ is associated with points in the same class. Secondly, we aim to evaluate whether the embedding space is also globally discriminant, i.e. that prototypes organize themselves to form clusters that match the classes, measured by two metrics. i) We associate with each prototype ${\mathbf{z}}_{k}$ a label using majority voting from the associated samples. Then we make use of the popular silhouette score (Rousseeuw, 1987) to evaluate whether prototypes’ positions are correctly aligned with these labels. ii) As commonly done to assess DR representations (Huang et al., 2022), we also consider the downstream task of clustering the prototypes using the k-means algorithm. We then assign to each $\{{\mathbf{x}}_{i}\}$ the cluster of its associated prototype, acting as a predicted label, and compute the Normalized Mutual Information (NMI) (Kvålseth, 2017) between these labels and $\{y_{i}\}$. More details about evaluation metrics can be found in Appendix F.
@@ 6 Conclusion
By making a connection between the GW problem and popular clustering and DR algorithms, we proposed a unifying framework denoted as DistR. DistR enables transforming any DR algorithm into a joint DR-clustering method that produces embeddings (or prototypes) with chosen granularity. We believe that the versatility of the GW framework will enable new extensions in unsupervised learning. For instance, the formalism associated with (semi-relaxed) GW barycenters naturally enables addressing multi-view settings with multiple unaligned inputs of different sizes. Other promising directions involve, better capturing multiple dependency scales in the input data by hierarchically adapting the resolution of the embedding similarity graph, or enabling batch optimization of embeddings to operate over much larger datasets.
Acknowledgments
The authors are grateful to Aurélien Garivier for insightful discussions. This project was supported in part by the ANR project OTTOPIA ANR-20-CHIA-0030 and through the MATTER project ANR-23-ERCC-0006-01. We gratefully acknowledge the support of the Centre Blaise Pascal’s IT test platform at ENS de Lyon (Lyon, France) for Machine Learning facilities. The platform operates the SIDUS solution (Quemener and Corvellec, 2013). This work benefited from state aid managed by the Agence Nationale de la Recherche under the France 2030 programme, reference ANR-23-IACL-0005. Finally, it received funding from the Fondation de l’École polytechnique and the Personalized Health and Related Technologies PHRT project 2022-644.
Appendix A Appendix
Outline of the appendix:
  • Appendix B: provide the proofs for the results stated in Section 3, namely the one for Lemma 3.1 in Section B.1; for Theorem 3.2 in Section B.2 and additional necessary and sufficient conditions under different assumptions as discussed in Remark 3.3 developed in Section B.3
  • Appendix C: provide the definition of weak isomorphism in the GW framework, proofs regarding the characterization of the generalized srGW discrepancy as a divergence, mentioned in Section 4.1 of the main paper.
  • Appendix D: proof of Theorem 4.1 on the clustering properties of srGW barycenters.
  • Appendix E: Algorithmic details for DistR under generic and low-rank settings.
  • Appendix F: Several additional information and results regarding the experiments detailed in Section 5 of the main paper.
    • Section F.1 provides details on methods’ implementation, validation of hyperparameters, datasets and metrics.
    • Section F.2 compares DistR with COOT clustering.
    • Section F.3: Best trade-off between metrics using t-SNE with $l_{2}$ symmetrization and UMAP.
    • Section F.4 reports complete scores on all datasets.
    • Section F.5 and F.6 study homogeneity vs silhouette and homogeneity vs kmeans NMI scores for various numbers of prototypes.
    • Section F.8 study the sensitivity of benchmarked methods to the embedding dimension $d$ using spectral DR methods.
    • Section F.9 compares computation time across methods.
    • Section F.10 detail our proofs of concepts with hyperbolic DR kernels.
Appendix B Proof of results in Section 3
B.1 Proof of lemma 3.1
@@ Appendix B Proof of results in Section 3
B.2 Proof of theorem 3.2
In the following $\operatorname{DS}$ is the space of $N\times N$ doubly stochastic matrices. We begin by proving the first point of theorem 3.2. We will rely on the simple, but useful, result below.
Proposition B.1.
@@
$\begin{split}\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij})&=\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},\sum_{kl}[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{kl}T_{ik}T_{jl})\\ &\leq\sum_{ijkl}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{kl})T_{ik}T_{jl}\,.\end{split}$ (7)
@@
$\min_{{\mathbf{Z}}}\min_{{\mathbf{T}}\in\operatorname{DS}}\ \sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij})\leq\min_{{\mathbf{Z}}}\ \min_{{\mathbf{T}}\in\operatorname{DS}}\sum_{ijkl}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{kl})T_{ik}T_{jl}\,.$ (8)
@@
For the first point it suffices to see that the condition eq. 11 implies that $\{{\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}:{\mathbf{Z}}\in\mathbb{R}^{N\times d},{\mathbf{T}}\in\operatorname{DS}\}\subseteq\{{\mathbf{C}}_{Z}({\mathbf{Z}}):{\mathbf{Z}}\in\mathbb{R}^{N\times d}\}$ (in fact we have equality by choosing ${\mathbf{T}}=\mathbf{I}_{N}$) and thus eq. 6 holds and we apply proposition B.1.
For the second point we will also show that eq. 6 holds. Consider ${\mathbf{Z}}^{\star},{\mathbf{T}}^{\star}$ minimizers of $\min_{{\mathbf{Z}},{\mathbf{T}}\in\operatorname{DS}}\ \sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij})$. By hypothesis there exists ${\mathbf{Y}}\in\mathbb{R}^{N\times d}$ such that $\forall(i,j)\in{[\![N]\!]}^{2},[{\mathbf{T}}^{\star}{\mathbf{C}}_{Z}({\mathbf{Z}}^{\star}){{\mathbf{T}}^{\star}}^{\top}]_{ij}\geq[{\mathbf{C}}_{Z}({\mathbf{Y}})]_{ij}$. Since $L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},\cdot)$ is non-decreasing for any $(i,j)$ then $\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}^{\star}{\mathbf{C}}_{Z}({\mathbf{Z}}^{\star}){{\mathbf{T}}^{\star}}^{\top}]_{ij})\geq\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Y}})]_{ij})$ and thus $\sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{T}}^{\star}{\mathbf{C}}_{Z}({\mathbf{Z}}^{\star}){{\mathbf{T}}^{\star}}^{\top}]_{ij})\geq\min_{{\mathbf{Z}}}\ \sum_{ij}L([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij})$ which gives the condition eq. 6 and we have the conclusion by proposition B.1. ∎
@@
    $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\sum_{ij}L_{2}([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},\langle{\mathbf{z}}_{i},{\mathbf{z}}_{j}\rangle)$ (13)
@@
For the first point we show that the condition eq. 11 of proposition B.2 is satisfied. Indeed take any ${\mathbf{Z}},{\mathbf{T}}$ then ${\mathbf{T}}{\mathbf{C}}_{Z}({\mathbf{Z}}){\mathbf{T}}^{\top}={\mathbf{T}}{\mathbf{Z}}{\mathbf{Z}}^{\top}{\mathbf{T}}^{\top}=({\mathbf{T}}{\mathbf{Z}})({\mathbf{T}}{\mathbf{Z}})^{\top}={\mathbf{C}}_{Z}({\mathbf{T}}{\mathbf{Z}})$.
For the second point if we consider $\tilde{L}(a,b)=a\times b$ then we use that the neighbor embedding problem eq. 16 is equivalent to $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\sum_{ij}-[{\mathbf{C}}_{X}({\mathbf{X}})]_{ij}\log([{\mathbf{C}}_{Z}({\mathbf{Z}})]_{ij})=\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\sum_{ij}\tilde{L}([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}})]_{ij})$ where $[\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}})]_{ij}=g({\mathbf{z}}_{i}-{\mathbf{z}}_{j})+\log(R({\mathbf{Z}}))$ with $g=-\log\circ f$. Since $f$ is logarithmically concave $g$ is convex. Moreover we have that $\tilde{L}(a,\cdot)$ is convex (it is linear) and non-decreasing since $a\in\mathbb{R}_{+}$ in this case (${\mathbf{C}}_{X}({\mathbf{X}})$ is non-negative). Also for any ${\mathbf{Z}}\in\mathbb{R}^{N\times d}$ and ${\mathbf{T}}\in\operatorname{DS}$ we have, using Jensen’s inequality since ${\mathbf{T}}$ is doubly-stochastic,
@@
Now we will prove that $\log(R({\mathbf{Z}}))\geq\log(R({\mathbf{T}}{\mathbf{Z}}))$. Using Birkhoff’s theorem (Birkhoff, 1946) the matrix ${\mathbf{T}}$ can be decomposed as a convex combination of permutation matrices, i.e. , ${\mathbf{T}}=\sum_{k}\lambda_{k}{\mathbf{P}}_{k}$ where $({\mathbf{P}}_{k})_{k}$ are permutation matrices and $\lambda_{k}\in\mathbb{R}_{+}$ with $\sum_{k}\lambda_{k}=1$. Hence by convexity and Jensen’s inequality $R({\mathbf{T}}{\mathbf{Z}})=R(\sum_{k}\lambda_{k}{\mathbf{P}}_{k}{\mathbf{Z}})\leq\sum_{k}\lambda_{k}R({\mathbf{P}}_{k}{\mathbf{Z}})$. Now using that $R$ is permutation invariant we get $R({\mathbf{P}}_{k}{\mathbf{Z}})=R({\mathbf{Z}})$ and thus $R({\mathbf{T}}{\mathbf{Z}})\leq\sum_{k}\lambda_{k}R({\mathbf{Z}})=R({\mathbf{Z}})$. Since the logarithm is non-decreasing we have $\log(R({\mathbf{Z}}))\geq\log(R({\mathbf{T}}{\mathbf{Z}}))$ and, overall,
@@
Thus if we introduce ${\mathbf{Y}}={\mathbf{T}}{\mathbf{Z}}$ we have $[{\mathbf{T}}\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}}){\mathbf{T}}^{\top}]_{ij}\geq[\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Y}})]_{ij}$ and eq. 12 is satisfied. Thus we can apply proposition B.2 and state that $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\sum_{ij}\tilde{L}([{\mathbf{C}}_{X}({\mathbf{X}})]_{ij},[\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}})]_{ij})$ and $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\operatorname{GW}_{\tilde{L}}({\mathbf{C}}_{X}({\mathbf{X}}),\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}}),\frac{1}{N}\mathbf{1}_{N},\frac{1}{N}\mathbf{1}_{N})$ are equivalent which concludes as $\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\operatorname{GW}_{\tilde{L}}({\mathbf{C}}_{X}({\mathbf{X}}),\widetilde{{\mathbf{C}}_{Z}}({\mathbf{Z}}),\frac{1}{N}\mathbf{1}_{N},\frac{1}{N}\mathbf{1}_{N})=\min_{{\mathbf{Z}}\in\mathbb{R}^{N\times d}}\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),\frac{1}{N}\mathbf{1}_{N},\frac{1}{N}\mathbf{1}_{N})$. ∎
@@
is actually concave. Indeed, in this case there exists a minimizer which is an extremal point of ${\mathcal{U}}(\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$. By Birkhoff’s theorem (Birkhoff, 1946) these extreme points are the matrices $\frac{1}{N}{\mathbf{P}}$ where ${\mathbf{P}}$ is a $N\times N$ permutation matrix. Consequently, when the function eq. 21 is concave minimizing $\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$ in ${\mathbf{Z}}$ is equivalent to minimizing in ${\mathbf{Z}}$
@@
where $a_{\mathbf{X}},b_{\mathbf{Z}}$ are terms that do not depend on ${\mathbf{T}}$. Since the problem is quadratic the concavity only depends on the term $-\sum_{ijkl}[{\mathbf{C}}_{X}({\mathbf{X}})]_{ik}\log([{\mathbf{K}}_{\mathbf{Z}}]_{jl})T_{ij}T_{kl}=-\operatorname{Tr}({\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}^{\top}\log({\mathbf{K}}_{\mathbf{Z}}){\mathbf{T}})$. From (Maron and Lipman, 2018) we know that the function ${\mathbf{T}}\to-\operatorname{Tr}({\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{T}}^{\top}\log({\mathbf{K}}_{\mathbf{Z}}){\mathbf{T}})$ is concave on ${\mathcal{U}}(\frac{1}{N}\bm{1}_{N},\frac{1}{N}\bm{1}_{N})$ when ${\mathbf{C}}_{X}({\mathbf{X}})$ is CPD and $\log({\mathbf{W}}_{\mathbf{Z}})$ is CPD. This concludes the proof.
B.3 Necessary and sufficient condition